Construct the confidence interval for the population variance for the given values. Round your answers to one decimal place. n=13 , s2=9.8, and c=0.98
Answer #1

s² = 9.8

n = 13

α = 1-0.98 = 0.02

Critical value, χ²α/2 = CHISQ.INV.RT(0.02/2, 12) = 26.2170

Critical value, χ²1-α/2 = CHISQ.INV.RT(1-0.02/2, 12) = 3.5706

98% Confidence interval for population variance :